UT 20130213 AO#2 SX

Observer:JCGuerra(LBTO),DMiller(LBTO)
Telescope Operator: David(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO)
Telescope Support: Jeff U(LBTO)
Instrument Support: Tim S (LBTO)

Summary

RR installed in right side (LFBG) issuetrack #4542

Detail

At the beginning of operation we found that we were no communicating from the slope computer to the AdSec. We checked fibers: all OK. We called Joe K: no changes. We powered off the WUint and Powered On and we had communication. The slope computer must not of initialized properly. Note, the fastlink test routine "aoidl test_fastlink.pro" returned fail even though the fastlink communication was working (we could see the frame counter on the AdSec Mirror GUI incrementing).

Finding spot: 2 hours

We started with the hexapod values from January list on the Adaptive Optics Twiki, procedures, RR. We performed several spiral searches with different centers and finally found a spot. We centered the right spot

Remove Coma: 1 hour

The return spot on the IRTC had big coma. We had set the IRTC to the small FOV and it was difficult to determine which was to move Rx,Y and Ry,X to remove coma. We went back to large FOV, used a low integration rate, moved the hexapod so out of focus and then was able to remove most of the coma. We then went to small FOV and touched up the coma. The hexapod parameters for a reasonable spot on the IRTC are:

Right Spot: x,y,z,rx,ry = -0.32, 1.7, 4.115, -380, -50

Left Spot: x,y,z,rx,ry = -0.32, 1.7, 4.06, -380, -66 (rough values, did not have time to remove coma)

Acquisition Camera Problems: 3 hours

Once we had the spot on nearly the correct spot on the IRTC, we did not see any light on the acquisition camera, CCD47. We moved the bayside stages and found the pyramid illuminated, but still no light on CCD47. We were able to close the loop and get an OK correction and decent image on the IRTC but did not collect any IRTC image (there was vibration and the loop was not stable so we could only use ~400 modes). We reset the BCU47 several times and even powered off the entire WUnit, but we never saw any light on CCD47. We shone a flashlight into the WUint, but CCD47 did not appear to be light sensitive. We even turned on the overhead lights, but not change.

Note, when we "Acuire TV dark", the CCD47 image does not change (lighter or darker) at quadrant boundaries, but maybe the upper left quadrant and 1/3 for the upper right quadrant will change the same amount at the same time. Then the rest of the upper right quadrant and 1/3 of the lower left quadrant will change, etc. Is this correct?

We created IT# 4547

-- JuanGuerra - 06 Mar 2013
Topic revision: r4 - 06 Mar 2013, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback